OpenStudy (anonymous):

michele Given the parent functions f(x) = log10 x and g(x) = 3x –1, what is f(x) ⋅g(x)?

OpenStudy (anonymous):

@Michele_Laino f(x) ⋅g(x) = log10 x –3x log10 x check to be right

OpenStudy (michele_laino):

I think that it is: \[\Large f\left( x \right) \cdot g\left( x \right) = \left( {{{\log }_{10}}x} \right)\left( {3x - 1} \right)\]

OpenStudy (anonymous):

A. f(x) ⋅g(x) = log10 (3x –1)^x B. f(x) ⋅g(x) = log10 x^(3x –1) C. f(x) • g(x) = 3x log10 x + log10 x D. f(x) ⋅g(x) = log10 x –3x log10 x @Michele_Laino which?

OpenStudy (anonymous):

which to be right michele?

OpenStudy (michele_laino):

yes! If we apply this property of logarithms: \[\Large m{\log _{10}}n = {\log _{10}}\left( {{n^m}} \right)\] with n=x, and m=3x-1, we get: \[\Large f\left( x \right) \cdot g\left( x \right) = \left( {{{\log }_{10}}x} \right)\left( {3x + 1} \right) = {\log _{10}}\left( {{x^{3x + 1}}} \right)\]

OpenStudy (anonymous):

is A answer?

OpenStudy (michele_laino):

A it is not, since we have 3x-1 as exponent, whereas in A we have x as exponent

OpenStudy (anonymous):

oh B is right!

OpenStudy (michele_laino):

that's right!